I need to know exact steps on how to put an inequality into slope intercept form. It would help to have a few examples.
y-2>2x+3
-2x+2y=-4

Found 3 solutions by Mathtut, Alan3354, solver91311:
Answer by Mathtut(3670)

slope intercept form is y=mx+b
:
y-2>2x+3
:
y>2x+5.......added 2 to each side
:
-2x+2y=-4
:
2y=2x-4....added 2x to both sides
:
y=x-2......divided all terms by 2

Answer by Alan3354(69443)

I need to know exact steps on how to put an inequality into slope intercept form. It would help to have a few examples.
----------------------------
It's simpler than it seems - you solve for y. That's all there is to it.
---------------
y-2>2x+3
y>2x + 5 That's it.
---------------
-2x+2y = -4
2y = 2x-4
y = x-2
We're done.

Answer by solver91311(24713)



In order to put any 2-variable equation or inequality into slope-intercept form, you simply need to solve the equation or inequality for y. What that means is that you must use the three basic operations: addition, multiplication, and raising to a power in an appropriate sequence so that the final result is y by itself on the left of the relationship sign and everything else in simplest form, i.e like terms collected, on the right. Remember though, when working with inequalities, if you multiply by a number less than zero, you must change the sense of the inequality.

Why didn't I mention subtraction, division, and taking a root? Simply because none of those things are truly basic operations. To subtract, you add the additive inverse, i.e. if you want to subtract 2, you add a -2. Division is multiplication by the reciprocal. If you want to divide by 2, you multiply by . And taking a root is the same as raising to a fractional power. Taking the square root is raising to the power.

For your first one:



Just add 2 to both sides:



For your second one, which is not an inequality but an equation:




Add 2x to both sides:



Multiply both sides by



John